Suppose a firm directly deposits its annual profit at one of the commercial banks in the country as a checkable deposit and that the reserve ratio is 10%. Trace out the effects of a $60 million profit of that firm on the value of checkable bank deposits. What is the size of the money multiplier?
1
Expert's answer
2020-09-28T08:02:12-0400

Checking Bank deposits will grow by 60 million if the firm transfers the entire amount of profit to the checking Deposit.

if checking deposits are included in the calculation of monetary aggregates, then the monetary multiplier is6

"m=\frac{1}{r}\times D=\frac{1}{0.1}\times60=600"


if checking deposits are not included in the calculation of monetary aggregates, then the monetary multiplier is:

"m=\frac{1}{r}=\frac{1}{0.1}=10"